OpenStudy (anonymous):

solve system by anymeathod 4x-7y=-13 3x-5y=-9

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Multiply both equations by 3 and -4, so you get \[12x-21y=-39\] and \[-12x+20y=36\] add to get \[-y=-3\] or \[y=3\] substitute and you get \[x=2\]

OpenStudy (anonymous):

long problem :/ especially to do online lol. So I, personally, would use substitution for this type of problem. 4x-7y=-13 3x-5y=-9 Second equation: y=3/5x+9/5 Substitute second equation to first equation. 4x-7(3/5x+9/5)=-13 4x-21/5x-63/5=-13 -1/5x=-2/5 x=2. substitute 2 for x into the second equation. 3(2)-5y=-9 6-5y=-9 -5y=-9/6 y=9/30 y=3/10. That should be right.